# Dwarf Galaxies in the MATLAS Survey: Hubble Space Telescope Observations of Nuclear Star Clusters

### v1

@@
e-mail: melina.poulain@oulu.fi, melina.poulain45@gmail.com 22institutetext: Institute für Astro- und Teilchenphysik, Universität Innsbruck, Technikerstraße 25/8, Innsbruck, A-6020, Austria 33institutetext: Université de Strasbourg, CNRS, Observatoire astronomique de Strasbourg, UMR 7550, F-67000 Strasbourg, France 44institutetext: UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh EH9 3HJ, UK 55institutetext: Dept. of Physics, Astronomy, Geology, and Environmental Sciences, Youngstown State University, Youngstown, OH 44555 USA 66institutetext: Department of Astronomy and Center for Galaxy Evolution Research, Yonsei University, Seoul 03722 77institutetext: INAF – Astronomical Observatory of Abruzzo, Via Maggini, 64100 Teramo, Italy 88institutetext: Institute of Physics, Laboratory of Astrophysics, École Polytechnique Fédéale de Lausanne (EPFL), 1290 Sauverny, Switzerland 99institutetext: Institute of Astronomy, Madingley Rd, Cambridge CB3 0HA, UK 1010institutetext: Visiting Fellow, Clare Hall, University of Cambridge, Cambridge, UK 1111institutetext: Department of Astronomy, Yonsei University, 50 Yonsei-ro Seodaemun-gu, Seoul, 03722, Republic of Korea 1212institutetext: Univ. Lyon, ENS de Lyon, Univ. Lyon 1, CNRS, Centre de Recherche Astrophysique de Lyon, UMR5574, 69007 Lyon, France
@@
In dwarf galaxies, nuclear star clusters (NSCs) are believed to primarily form from the migration and merger of globular clusters (GCs), with a possible contribution from in-situ star-forming activity triggered by gas infall. We present the study of NSCs in 41 MATLAS survey dwarf galaxies including ultra-diffuse galaxies (UDGs), as part of a large follow-up imaging program with the Hubble Space Telescope (HST) Advanced Camera for Surveys (ACS) using the F606W and F814W filters. The sample is biased towards low-surface brightness and large dwarfs, i.e., UDG-like galaxies, and includes two galaxies with a double nucleus, 13 newly identified nucleated dwarfs thanks to HST’s high spatial resolution, and five candidate ultra-compact dwarf progenitors. We modeled the NSCs with a Sérsic profile and derived their structural properties and photometry. We find the NSC Sérsic index to increase with the luminosity and stellar mass, while no obvious trend is seen on the effective radius and ellipticity. The faint NSCs tend to have a constant color profile, whereas the bright ones have a bluer center, suggesting that the most massive NSCs in our sample might have experienced a mixed formation scenario, including in-situ star formation. A significant portion of our NSCs tend to be more massive than for other galaxy samples of similar stellar mass, which could be due to some dwarfs ongoing tidal disruption or an initial formation of massive NSCs from multiple GC mergers and in-situ star forming activity. More observations of resolved NSC are needed to be able to infer their formation scenario from the structural properties and photometry in dwarfs.
@@
Nuclear star clusters (NSCs) are known to be the densest stellar systems in the Universe. These small compact objects usually have an effective radius within ten parsecs and a stellar mass M from to M⊙ (Neumayer et al. 2020). Although they reside in the central region of galaxies, their presence depends on the stellar mass of the host (M). This so-called nucleation fraction reaches a maximum of more than for galaxies with M109 M⊙ (van den Bergh 1986; Ferguson & Binggeli 1994; den Brok et al. 2014; Ordenes-Briceño et al. 2018; Sánchez-Janssen et al. 2019), while an absence of NSC is observed for more than 90% of dwarf and high-mass galaxies with M M⊙ and M M⊙, respectively (e.g., Sandage et al. 1985; Faber et al. 1997; Côté et al. 2006; Antonini et al. 2015; Poulain et al. 2021; Carlsten et al. 2022; Su et al. 2022; Marleau et al. 2025).
@@
Two main pathways appear to take part in the formation of NSCs. The first one is the migration scenario, where GCs form in a galaxy, migrate towards its center due to dynamical friction, and finally merge to form a NSC (Tremaine et al. 1975; Lotz et al. 2001; Bekki 2007). This scenario has been reproduced well by N-body simulations (e.g., Antonini et al. 2012; Lyubenova & Tsatsi 2019), and can explain the formation of NSCs with old metal-poor stellar populations. The second one is the in-situ scenario (Loose et al. 1982), characterized by the formation of NSCs from star-forming activity caused by infalling gas. While the resulting stellar population tends to be young and metal-rich, the in-situ scenario is not the only way to produce such young NSCs. In Milky-Way-like galaxies, N-body simulations have shown that NSCs can also be formed by the merging of young, metal-rich clusters (Arca-Sedda et al. 2015; Schiavi et al. 2021). The variety of stellar populations of some NSCs, like M54 in the Sagittarius dwarf, imply a formation based on both the migration and in-situ processes (e.g., Sills et al. 2019; Alfaro-Cuello et al. 2019; Kacharov et al. 2022). An example of mixed scenario is the wet migration scenario (Guillard et al. 2016), based on hydrodynamical simulations in dwarf galaxies, which proposes the formation of a GC-like star cluster in a gas-rich dwarf which contains a gas reservoir allowing it to maintain star-forming activity while migrating towards the galaxy center, producing a rather young, blue NSC. Several recent studies have looked for the main formation channel of NSCs and their results suggest that the migration scenario dominates in galaxies with M¡ M⊙, while the in-situ formation prevails in more massive galaxies, and both seem to contribute equally around M⊙ (e.g., Turner et al. 2012; den Brok et al. 2014; Johnston et al. 2020; Fahrion et al. 2021, 2022a, 2022b, 2024).
@@
Together with theoretical works, the formation of NSCs is traditionally studied using three main types of observational data. The first are deep imaging surveys, undertaken mostly via ground-based observations. They are very efficient to unveil large numbers of new NSCs using photometric data and offer detailed statistics about the morphology and local environment of the nucleated galaxies (e.g., Reaves 1983; Binggeli et al. 1985; Ordenes-Briceño et al. 2018; Sánchez-Janssen et al. 2019; Habas et al. 2020; Carlsten et al. 2022; Su et al. 2022; Paudel et al. 2023). The second is the analysis of the stellar populations using spectrographs. The Space Telescope Imaging Spectrograph (STIS) of the Hubble Space Telescope (HST) was used by Rossa et al. (2006) to produce the largest spectroscopic survey of NSCs in late-type galaxies (LTGs) to date. Lately, the Multi Unit Spectroscopic Explorer (MUSE) and FOcal Reducer and low dispersion Spectrograph (FORS) instruments on the Very Large telescope (VLT) were favored to investigate the age, metallicity, and star formation history of NSCs in dwarf galaxies (Paudel et al. 2011; Johnston et al. 2020; Fahrion et al. 2021, 2022a). Finally, the third is the study of their structural properties with high-resolution space-based observations. HST has been frequently used throughout the years for this kind of work. The luminosity and size of NSCs were first measured in LTGs (see for example Carollo et al. 1997; Matthews et al. 1999; Böker et al. 2002; Balcells et al. 2003; Böker et al. 2004; Balcells et al. 2007; Georgiev & Böker 2014), before targeting early-type galaxies (ETGs) with the HST surveys in the Virgo, Fornax and Coma clusters (Lotz et al. 2001; Graham & Guzmán 2003; Côté et al. 2006; Turner et al. 2012; den Brok et al. 2014) where more than 200 NSCs were characterized. However those studies are lacking statistics for dwarf galaxies with M¡ M⊙. More recently, the Early Release Observations from the space telescope Euclid (Marleau et al. 2025) and additional studies with HST (e.g., Spengler et al. 2017; Pechetti et al. 2020; Zanatta et al. 2021; Hoyer et al. 2023) have highlighted new NSCs in galaxy clusters and within the Local Volume (LV), especially in dwarf galaxies.
@@
Related to NSCs are the GCs which also populate the halos of galaxies. Comparing the GCs to the NSCs properties is important in the context of the migration scenario, especially in dwarf galaxies where this formation path is favored due to their shorter dynamical friction timescales (Hernandez & Gilmore 1998; Oh & Lin 2000). Several studies already pointed out correlations between these two types of star clusters. In terms of spatial distribution, a lack of bright GCs was observed in the center of nucleated elliptical dwarfs in the Virgo and Fornax clusters by Lotz et al. (2001). A deficit of GCs at small radii in both nucleated and non-nucleated UDGs was reported by Marleau et al. (2024), and could be caused by either dynamical friction or the stalling of GCs at the core radius in a core dark matter halo. For dwarf galaxies located in galaxy clusters and in the Local Volume (LV), NSCs and GCs also seem to share a similar occupation fraction, dependant on the stellar mass of the host galaxy (Sánchez-Janssen et al. 2019; Carlsten et al. 2022).
@@
Beyond the formation of NSC, it is also interesting to look for nucleated galaxies ongoing a transition into ultra compact dwarfs galaxies (UCDs). UCDs (Hilker et al. 1999; Drinkwater et al. 2000; Phillipps et al. 2001) are compact stellar systems with an effective radius below 100 pc and a stellar mass above M⊙ (Evstigneeva et al. 2007; Mieske et al. 2008; Misgeld & Hilker 2011). UCDs have been reported in a wide range of environments, from galaxy clusters, galaxy groups, and around field galaxies (e.g., Madrid et al. 2010; Chilingarian et al. 2011; Norris & Kannappan 2011; Faifer et al. 2017; De Bórtoli et al. 2020; Liu et al. 2020; Saifollahi et al. 2021). In the past two decades, a number of scenarios have been put forward to explain the nature of UCDs, and these can broadly be placed into two different formation channels. A fraction of the UCDs are expected to correspond to the high-mass end of the GC population (Kroupa 1998; Mieske et al. 2002, 2012), while the most massive UCDs are likely to be NSC remnants of tidally stripped nucleated dwarf galaxies (e.g., Bassino et al. 1994; Bekki et al. 2001; Goerdt et al. 2008; Mayes et al. 2021; Wang et al. 2023).
@@
Studying dwarf galaxies is important to comprehend the formation and evolution of NSCs. However, only few observing programs using high spatial resolution imaging exist, and most of them focused galaxy clusters. The sample of 2210 dwarf galaxies identified in the Mass Assembly of early-Type GaLAxies with their fine Structures (MATLAS) survey offers large statistics beyond the LV, up to a distance of about 45 Mpc, and outside the environment of galaxy clusters (Habas et al. 2020). A photometric study based on the MATLAS deep images from the Canada-France-Hawaii telescope (CFHT) was already undertaken on this sample, unveiling 507 nucleated dwarfs (Poulain et al. 2021). HST observations were conducted for 79 dwarf and ultra-diffuse galaxies (UDGs) from the MATLAS survey (Marleau et al. 2024), 41 of them being nucleated. UDGs are defined as dwarf galaxies with an unusually large effective radius and low central surface brightness, typically Re¿1.5 kpc and ¿24 mag/arcsec2, respectively (van Dokkum et al. 2015). Designed to characterize their GC populations, these high resolution observations marginally resolve the nuclear region of the galaxies and allow for a more detailed study of the NSCs than previously performed in Poulain et al. (2021). A first paper on the NSCs by Poulain et al. (2025) focused on the detection of multiple star clusters and stellar tidal tails in a handful of dwarfs. The observed tidal tails were successfully reproduced in N-body simulations, and suggest that they are produced by ongoing star clusters mergers. In this paper we derive the NSCs photometric and structural properties in the 41 dwarf galaxies and UDGs, and discuss their possible formation and evolution pathways.
@@
This work is organized as follows. We introduce the MATLAS survey, as well as the dwarf galaxy sample in Section 2. We present the HST observations in Section 3, and describe the sample of nucleated dwarf galaxies and UDGs in Section 4. The method used to derive the properties of the NSCs are detailed in Section 5 and their structural and photometric properties are shown in Section 6. We discuss the results in terms of formation of NSCs in dwarf galaxies in Section 7, and their possible and evolution into UCDs in Section 8. Finally, we summarize and draw conclusions in Section 9.
@@ 2 The MATLAS dwarf sample
The galaxies selected for this work were identified in the MATLAS111http://matlas.astro.unistra.fr/WP/ survey (Duc et al. 2015; Duc 2020; Bílek et al. 2020). This survey was designed to study the outermost stellar populations, low surface brightness structures, globular clusters, and dwarf galaxy satellites of ETGs in the context of their mass assembly and the build-up of their scaling relations. The observations were carried out with MegaCam on the CFHT from 2012 to 2015. After being processed by the Elixir-LSB pipeline (Magnier & Cuillandre 2004), the resulting images are suited for the detection and study of low surface brightness structures, with a surface brightness limit down to mag/arcsec2 in the -band. The ETGs were selected from the sample of the ATLAS3D project (Cappellari et al. 2011), which is composed of multiwalength observations of a complete sample of 260 ETGs, numerical simulations, and semi-analytic modeling of galaxy formation. The 260 ETGs are located within a 45 Mpc volume and beyond the LV, with Galactic declinations , Galactic latitudes above 15°, and absolute magnitudes in the -band brighter than . The MATLAS sample is composed of 180 ETGs outside the environment of galaxy clusters, thus avoiding the Virgo cluster ETGs imaged in the Next Generation Virgo Survey (NGVS; Ferrarese et al. 2012). We note that this sample also contains 59 late-type galaxies (LTGs) following the same selection criteria as the ETGs, aside from the morphology. The data is divided into 150, 148, and 78 fields in the , , and -band, respectively, while the 12 closest ETGs were also observed in the -band. The MATLAS images have a field-of-view of and an average seeing of 0.96″, 0.84″, 0.65″, and 1.03″ in the , , , and -band, respectively.
@@
The -band fields were chosen for the detection and identification of dwarf galaxies due to their more complete sky coverage as compared to the other bands. The selection of the dwarfs (see Habas et al. 2020 for more details) was based on a visual inspection of the images coupled with a semi-automated method making use of source extractor (Bertin & Arnouts 1996). The final sample contains 2210 galaxies and includes 73% of elliptical dwarfs (dEs), 27% of irregular dwarfs (dIs), as determined from the CFHT images. Moreover, 23% of the dwarfs were classified as nucleated. From the spectroscopic redshifts available in the literature and derived heliocentric velocities from HI measurements, Habas et al. (2020) found that about 90% of the dwarfs with known distance, i.e., 14% of the sample, are satellites of one of the ETGs or LTGs from ATLAS3D, with 80% being satellites of the central ETGs in the MATLAS fields. Based on these results, it is reasonable to assume that a dwarf without a known distance is a satellite of the targeted ETG in the field it was identified and thus is located at the same distance. A recent spectroscopic study of 56 MATLAS dwarfs observed with the MUSE instrument on the VLT allowed us to obtain 52 new distances from the derived recessional velocities of the galaxies (Heesters et al. 2023), 75% of which – and 79% of the dEs – are consistent with the distance of the assumed host ETG. Therefore, for the remainder of this paper, we will make use of the measured distances of the dwarfs, when available, or fix the distance to that of the targeted ETG otherwise.
@@
The photometric and structural properties of the MATLAS dwarfs, and that of their nuclei, were studied in Poulain et al. (2021). Follow-up HST observations have been performed to study the GC population of selected UDGs from Marleau et al. (2021). The first observations focused on MATLAS-2019, which has recently become an object of interest due to its large GC population (Müller et al. 2021; Danieli et al. 2022). At least 11 of the GCs have been confirmed by MUSE spectroscopy (Müller et al. 2020), strengthening the capability of detecting star clusters with HST and select them based on their color and concentration index. A more recent work made use of a similar method to investigate the GC population of 74 UDGs (Marleau et al. 2024). These follow-up observations are well suited for a detailed analysis of the NSCs of the nucleated galaxies.

### v2

@@
e-mail: melina.poulain@oulu.fi, melina.poulain45@gmail.com 22institutetext: Institute für Astro- und Teilchenphysik, Universität Innsbruck, Technikerstraße 25/8, Innsbruck, A-6020, Austria 33institutetext: Université de Strasbourg, CNRS, Observatoire astronomique de Strasbourg, UMR 7550, F-67000 Strasbourg, France 44institutetext: UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh EH9 3HJ, UK 55institutetext: Dept. of Physics, Astronomy, Geology, and Environmental Sciences, Youngstown State University, Youngstown, OH 44555 USA 66institutetext: Department of Astronomy and Center for Galaxy Evolution Research, Yonsei University, Seoul 03722 77institutetext: INAF – Astronomical Observatory of Abruzzo, Via Maggini, 64100 Teramo, Italy 88institutetext: Institute of Physics, Laboratory of Astrophysics, École Polytechnique Fédéale de Lausanne (EPFL), 1290 Sauverny, Switzerland 99institutetext: Institute of Astronomy, Madingley Rd, Cambridge CB3 0HA, UK 1010institutetext: Visiting Fellow, Clare Hall, University of Cambridge, Cambridge, UK 1111institutetext: Division of Science Education, Kangwon National University, Chuncheon, Republic of Korea 1212institutetext: Department of Astronomy, Yonsei University, 50 Yonsei-ro Seodaemun-gu, Seoul, 03722, Republic of Korea 1313institutetext: Univ. Lyon, ENS de Lyon, Univ. Lyon 1, CNRS, Centre de Recherche Astrophysique de Lyon, UMR5574, 69007 Lyon, France
@@
In dwarf galaxies, nuclear star clusters (NSCs) are believed to primarily form from the migration and merger of globular clusters (GCs), with a possible contribution from in situ star-forming activity triggered by gas infall. We present the study of NSCs in 41 MATLAS survey dwarf galaxies including ultra-diffuse galaxies (UDGs), as part of a large follow-up imaging program with the Hubble Space Telescope (HST) Advanced Camera for Surveys (ACS) using the F606W and F814W filters. The sample is biased toward low surface brightness and large dwarfs, i.e., UDG-like galaxies, and includes two galaxies with a double nucleus; 13 newly identified nucleated dwarfs, thanks to HST’s high spatial resolution; and five candidate ultra-compact dwarf progenitors. We modeled the NSCs with a Sérsic profile and derived their structural properties and photometry. We find the NSC Sérsic index to increase with the luminosity and stellar mass, while no obvious trend is seen for the effective radius and ellipticity. The faint NSCs tend to have a constant color profile, whereas the bright ones have a bluer center, suggesting that the most massive NSCs in our sample might have experienced a mixed formation scenario, including in situ star formation. A significant portion of our NSCs tend to be more massive than for other galaxy samples of similar stellar mass, which could be due to some dwarfs undergoing tidal disruption or an initial formation of massive NSCs from multiple GC mergers and in situ star-forming activity. More observations of resolved NSCs are needed to be able to infer their formation scenario from the structural properties and photometry in dwarfs.
@@
Nuclear star clusters (NSCs) are known to be the densest stellar systems in the Universe. These small compact objects usually have an effective radius within 10 parsec and a stellar mass M from to M⊙ (Neumayer2020). Although they reside in the central region of galaxies, their presence depends on the stellar mass of the host (M). Known as the nucleation fraction, it reaches a maximum of more than for galaxies with M109 M⊙ (VandenBergh1986; Ferguson1994; denBrok2014; Ordenes2018; Janssen2019), while an absence of NSC is observed for more than 90% of dwarf and high-mass galaxies with M M⊙ and M M⊙, respectively (e.g., Sandage1985; Faber1997; Cote2006; Antonini2015; Poulain2021; Carlsten2022; Su2022; Marleau2024a).
@@
Two main pathways appear to take part in the formation of NSCs. The first is the migration scenario, where GCs form in a galaxy, migrate toward its center due to dynamical friction, and finally merge to form a NSC (Tremaine1975; lotz2001; Bekki2007). This scenario has been reproduced well by N-body simulations (e.g., Antonini2012; Lyubenova2019), and can explain the formation of NSCs with old metal-poor stellar populations. The second pathway is the in situ scenario (Loose1982), characterized by the formation of NSCs from star-forming activity caused by infalling gas. While the resulting stellar population tends to be young and metal rich, the in situ scenario is not the only way to produce such young NSCs. In Milky Way (MW)-like galaxies, N-body simulations have shown that NSCs can also be formed by the merging of young metal-rich clusters (Arca-Sedda2015; Schiavi2021). The variety of stellar populations of some NSCs, for example M54 in the Sagittarius dwarf, imply a formation based on the migration and on in situ processes (e.g., Sills2019; AlfaroCuello2019; Kacharov2022). An example of a mixed scenario is the wet migration scenario (Guillard2016), based on hydrodynamical simulations in dwarf galaxies. This scenario proposes the formation of a GC-like star cluster in a gas-rich dwarf galaxy that contains a gas reservoir. The reservoir allows the cluster to maintain star-forming activity while it migrates toward the galaxy center, producing a rather young blue NSC. Several recent studies have looked for the main formation channel of NSCs and their results suggest that the migration scenario dominates in galaxies with M¡ M⊙, while the in situ formation prevails in more massive galaxies, and both seem to contribute equally around M⊙ (e.g., Turner2012; denBrok2014; Johnston2020; Fahrion2021; Fahrion2022b; Fahrion2022; Fahrion2024).
@@
Together with theoretical works, the formation of NSCs is traditionally studied using three main types of observational data. The first are deep imaging surveys, undertaken mostly via ground-based observations. They are very efficient at unveiling large numbers of new NSCs using photometric data, and they offer detailed statistics about the morphology and local environment of the nucleated galaxies (e.g., Reaves1983; Binggeli1985; Ordenes2018; Janssen2019; Habas2020; Carlsten2022; Su2022; Paudel2023). The second is the analysis of stellar populations using spectrographs. The Space Telescope Imaging Spectrograph (STIS) of the Hubble Space Telescope (HST) was used by Rossa2006 to produce the largest spectroscopic survey of NSCs in late-type galaxies (LTGs) to date. Recently, the Multi Unit Spectroscopic Explorer (MUSE) and FOcal Reducer and low dispersion Spectrograph (FORS) instruments on the Very Large telescope (VLT) were chosen to investigate the age, metallicity, and star formation history of NSCs in dwarf galaxies (Paudel2011; Johnston2020; Fahrion2021; Fahrion2022b). Finally, the third is the study of their structural properties with high-resolution space-based observations. HST has frequently been used throughout the years for this kind of work. The luminosity and size of NSCs were first measured in LTGs (see, e.g., Carollo1997; Matthew1999; Boeker2002; Balcells2003; Boeker2004; Balcells2007; Georgiev2014), before targeting early-type galaxies (ETGs) with the HST surveys in the Virgo, Fornax, and Coma clusters (lotz2001; Graham2003; Cote2006; Turner2012; denBrok2014) where more than 200 NSCs were characterized. However, those studies are lacking statistics for dwarf galaxies with M¡ M⊙. More recently, the Early Release Observations from the space telescope Euclid (Marleau2024a) and additional studies with HST (e.g., Spengler2017; Pechetti2020; Zanatta2021; Hoyer2023) have highlighted new NSCs in galaxy clusters and within the Local Volume (LV), especially in dwarf galaxies.
@@
Related to NSCs are the GCs that also populate the halos of galaxies. Comparing the properties of the GCs to those of the NSCs is important in the context of the migration scenario, especially in dwarf galaxies where this formation path is favored due to their shorter dynamical friction timescales (Hernandez1998; OhLin2000). Several studies already pointed out correlations between these two types of star clusters. In terms of spatial distribution, a lack of bright GCs was observed in the center of nucleated elliptical dwarfs in the Virgo and Fornax clusters by lotz2001. A deficit of GCs at small radii in both nucleated and non-nucleated UDGs was reported by Marleau2024, and could be caused by either dynamical friction or the stalling of GCs at the core radius in a core dark matter halo. For dwarf galaxies located in galaxy clusters and in the Local Volume (LV), NSCs and GCs also seem to share a similar occupation fraction, dependant on the stellar mass of the host galaxy (Janssen2019; Carlsten2022).
@@
Beyond the formation of NSC, it is also interesting to look for nucleated galaxies undergoing a transition into ultra-compact dwarf galaxies (UCDs). UCDs (Hilker1999; Drinkwater2000; Phillipps2001) are compact stellar systems with an effective radius below 100 pc and a stellar mass above M⊙ (Evstigneeva2007; Mieske2008; Misgeld2011). UCDs have been reported in a wide range of environments, in galaxy clusters, galaxy groups, and around field galaxies (e.g., Madrid2010; Chilingarian2011; Norris2011; Faifer2017; Bortoli2020; Liu2020; Saifollahi2021). In the past two decades a number of scenarios have been put forward to explain the nature of UCDs, and these can broadly be placed into two different formation channels. A fraction of the UCDs are expected to correspond to the high-mass end of the GC population (Kroupa1998; Mieske2002; Mieske2012), while the most massive UCDs are likely to be NSC remnants of tidally stripped nucleated dwarf galaxies (e.g., Bassino1994; Bekki2001; Goerdt2008; Mayes2021; Wang2023).
@@
Studying dwarf galaxies is important to comprehend the formation and evolution of NSCs. However, only a few observing programs using high spatial resolution imaging exist; most of them focus on galaxy clusters. The sample of 2210 dwarf galaxies identified in the Mass Assembly of early-Type GaLAxies with their fine Structures (MATLAS) survey offers large statistics beyond the LV, up to a distance of about 45 Mpc, and outside the environment of galaxy clusters (Habas2020). A photometric study based on the MATLAS deep images from the Canada-France-Hawaii telescope (CFHT) was already undertaken on this sample, and revealed 507 nucleated dwarfs (Poulain2021). HST observations were conducted for 79 dwarf and ultra-diffuse galaxies (UDGs) from the MATLAS survey (Marleau2024), 41 of which were nucleated. UDGs are defined as dwarf galaxies with an unusually large effective radius and low central surface brightness, typically Re¿1.5 kpc and ¿24 mag/arcsec2, respectively (vanDokkum2015). Designed to characterize their GC populations, these high-resolution observations marginally resolve the nuclear region of the galaxies and allow a more detailed study of the NSCs than previously performed in Poulain2021. A first paper on the NSCs by Poulain2025 focused on the detection of multiple star clusters and stellar tidal tails in a handful of dwarfs. The observed tidal tails were successfully reproduced in N-body simulations, and suggest that they are produced by ongoing star cluster mergers. In this paper we derive the NSCs photometric and structural properties in the 41 dwarf galaxies and UDGs, and discuss their possible formation and evolution pathways.
@@
This work is organized as follows. We introduce the MATLAS survey, as well as the dwarf galaxy sample in Sect. 2. We present the HST observations in Sect. 3, and describe the sample of nucleated dwarf galaxies and UDGs in Sect. 4. The method used to derive the properties of the NSCs is detailed in Sect. 5 and their structural and photometric properties are shown in Sect. 6. We discuss the results in terms of formation of NSCs in dwarf galaxies in Sect. 7, and their possible and evolution into UCDs in Sect. 8. Finally, we summarize and draw our conclusions in Sect. 9.
@@ 2 The MATLAS dwarf sample
The galaxies selected for this work were identified in the MATLAS111http://matlas.astro.unistra.fr/WP/ survey (Duc2015; Duc2020; Bilek2020). This survey was designed to study the outermost stellar populations, low surface brightness structures, globular clusters, and dwarf galaxy satellites of ETGs in the context of their mass assembly and the build-up of their scaling relations. The observations were carried out with MegaCam on the CFHT from 2012 to 2015. After being processed by the Elixir-LSB pipeline (MagnierCuillandre2004), the resulting images are suited for the detection and study of low surface brightness structures, with a surface brightness limit down to mag/arcsec2 in the -band. The ETGs were selected from the sample of the ATLAS3D project (Cappellari2011), which is composed of multiwavelength observations of a complete sample of 260 ETGs, numerical simulations, and semi-analytic modeling of galaxy formation. The 260 ETGs are located within a 45 Mpc volume and beyond the LV, with Galactic declinations , Galactic latitudes above 15°, and absolute magnitudes in the -band brighter than . The MATLAS sample is composed of 180 ETGs outside the environment of galaxy clusters, thus avoiding the Virgo cluster ETGs imaged in the Next Generation Virgo Survey (NGVS; Ferrarese2012). We note that this sample also contains 59 late-type galaxies (LTGs) following the same selection criteria as the ETGs, aside from the morphology. The data is divided into 150, 148, and 78 fields in the -, -, and -band, respectively, while the 12 closest ETGs were also observed in the -band. The MATLAS images have a field of view of and an average seeing of 0.96″, 0.84″, 0.65″, and 1.03″ in the -, -, -, and -band, respectively.
@@
The -band fields were chosen for the detection and identification of dwarf galaxies, due to their more complete sky coverage as compared to the other bands. The selection of the dwarfs (see Habas2020 for more details) was based on a visual inspection of the images coupled with a semi-automated method making use of source extractor (Bertin1996). The final sample contains 2210 galaxies and includes 73% of elliptical dwarfs (dEs), 27% of irregular dwarfs (dIs), as determined from the CFHT images. Moreover, 23% of the dwarfs were classified as nucleated. From the spectroscopic redshifts available in the literature and derived heliocentric velocities from HI measurements, Habas2020 found that about 90% of the dwarfs with known distance, i.e., 14% of the sample, are satellites of one of the ETGs or LTGs from ATLAS3D; 80% are satellites of the central ETGs in the MATLAS fields. Based on these results, it is reasonable to assume that a dwarf without a known distance is a satellite of the targeted ETG in the field it was identified, and thus is located at the same distance. A recent spectroscopic study of 56 MATLAS dwarfs observed with the MUSE instrument on the VLT allowed us to obtain 52 new distances from the derived recessional velocities of the galaxies (heesters2023), 75% of which (and 79% of the dEs) are consistent with the distance of the assumed host ETG. Therefore, for the remainder of this paper, we make use of the measured distances of the dwarfs, when available, or fix the distance to that of the targeted ETG otherwise.
@@
The photometric and structural properties of the MATLAS dwarfs, and that of their nuclei, were studied in Poulain2021. Follow-up HST observations were performed to study the GC population of selected UDGs from Marleau2021. The first observations focused on MATLAS-2019, which has recently become an object of interest due to its large GC population (Mueller2021; Danieli2022). At least 11 of the GCs have been confirmed by MUSE spectroscopy (Mueller2020), strengthening the capability of detecting star clusters with HST and selecting them based on their color and concentration index. A more recent work made use of a similar method to investigate the GC population of 74 UDGs (Marleau2024). These follow-up observations are well suited for a detailed analysis of the NSCs of the nucleated galaxies.